The blood platelet counts of a group of women have a​ bell-shaped distribution with a mean of

264.6264.6

and a standard deviation of

62.562.5.

​(All units are 1000

​cells/muμ​L.)

Using the empirical​ rule, find each approximate percentage below.

a.

What is the approximate percentage of women with platelet counts within

33

standard

deviationsdeviations

of the​ mean, or between

77.177.1

and

452.1452.1​?

b.

What is the approximate percentage of women with platelet counts between

139.6139.6

and

389.6389.6​?

a. Approximately

nothing​%

of women in this group have platelet counts within

33

standard

deviationsdeviations

of the​ mean, or between

77.177.1

and

452.1452.1.
Answer #1

Given \mu = 264.6 and \sigma =77.1

a) Z_{1} =\frac{X_{1}-\mu }{\sigma }=\frac{77.1-264.6}{62.5}=\frac{-187.5}{62.5}=-3

Z_{2 }=\frac{X_{2}-\mu }{\sigma }=\frac{452.1-264.6}{62.5}=\frac{187.5}{62.5}=3

P(77.1< X< 452.1)=P(-3< Z< 3)=99.7\%

b) Z_{1} =\frac{X_{1}-\mu }{\sigma }=\frac{139.6-264.6}{62.5}=\frac{-125}{62.5}=-2

Z_{2} =\frac{X_{2}-\mu }{\sigma }=\frac{389.6-264.6}{62.5}=\frac{125}{62.5}=2

P(139.6< X< 389.6)=P(-2 < Z< 2)=95 \%